# Gauge Symmetries, Contact Reduction, and Singular Field Theories

### v1

@@ X.2 The Pre-multicontact Constraint Algorithm
which is clearly an element of $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$. Upon substituting this into the dynamical equations $\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0$ and $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$, we must ensure that the resulting expressions do not contain any inconsistencies. Supposing that we do in fact obtain a compatible set of equations, we now restrict $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (96); in general, this step gives rise to further consistency conditions, and possibly to new constraints.
@@ XI Contact Reduction of Singular Field Theories
We shall now present the field-theoretic construction, which generalises the concepts developed in section (VII); as stated previously, we shall favour the Lagrangian formalism, as opposed to the Hamiltonian framework, as the latter suffers from significantly more pathological features, when considering singular field theories. Our exposition follows closely that of [29], and so we shall focus on features of the reduction process that are unique to singular theories.
Referring to the notation developed throughout section (IX.1), we suppose that $(J^{1}E,\Omega_{L})$ is a pre-multisymplectic Lagrangian system, with corresponding Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$. From the Cartan $d$-form $\Theta_{L}$, we define
@@
The vector field $\Sigma\in\mathfrak{X}^{\infty}(J^{1}E)$ then constitutes a scaling symmetry of degree $\Lambda$ if
$\mathfrak{L}_{\Sigma}L=\Lambda L\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}\quad\textrm{for all}\;\mu=0,\,\cdots,d-1$
@@
It is relatively straightforward to show that the equations of motion of $L$ directly imply the Herglotz-Lagrange field equations for $L^{H}$, when the former is restricted to the subspace upon which $L^{H}$ is defined. In light of this, suppose that, having identified a scaling symmetry $\Sigma$ of our pre-multisymplectic system, we adopt coordinates on $J^{1}E$ in such a way so as to render this vector field of the form
@@
The Euler-Lagrange field equation for $\rho$ then implies that the function $f$ may be written as
@@
which, when compared to (98), and recalling that the Herglotz-Lagrange equations require that $\partial_{\mu}s^{\mu}=L^{H}$, suggests we should identify
@@
All steps outlined above are applicable to regular Lagrangian systems; however, the novelty that arises in the singular case, is that the Lagrangian $d$-form $\Theta_{L^{H}}$ calculated from the symmetry-reduced function $L^{H}$ is not guaranteed to be pre-multicontact. Consequently, as a crucial additional step, we must compute the characteristic and Reeb distributions, after making the ‘naïve’ contact reduction; if these objects do not satisfy the conditions given in section (X), the symmetry reduction procedure fails. Of course, we shall only consider those physically interesting cases, in which the reduced space does inherit a pre-multicontact structure; nevertheless, it is a non-trivial condition that must be verified.
Having calculated $\Theta_{L^{H}}$, and confirmed that $\textrm{rank}\,\mathcal{D}^{R}=d+k$ and $\textrm{rank}\,\mathcal{C}=k$, for some $0<k\leqslant n(1+d)$, we have a pre-multicontact manifold $(\mathcal{S}_{0},\Theta_{L^{H}})$, in which $\mathcal{S}_{0}$ denotes the reduced space, obtained as a result of excising the scaling variable $\rho$. Geometrically speaking, if the system of interest is such that we may globally choose a single covariant configuration space variable as a scaling function: that is to say, if the quantity $\xi$ in (100) corresponds to a single field variable on $E$, then, upon writing the Lagrangian in the form (101), $E$ separates into two connected pieces $E_{\pm}$, where each of $E_{\pm}$ is the product of a trivial bundle, and a codimension-1 subspace: $E_{\pm}\cong\widetilde{E}\times_{M}(M\times\mathbb{R}_{\pm})$. The two components $E_{\pm}$ arise as a result of the change of variable $\xi\rightarrow e^{\rho/\Lambda}$, which selects only the positive component of $\xi$. In order to cover the full range of $\xi$, we must also consider a change of variables in which $\xi\rightarrow-\,e^{\rho/\Lambda}$. Within each of $E_{\pm}$, it is precisely the trivial bundle $M\times\mathbb{R}_{\pm}\rightarrow M$ that is eliminated by the contact reduction. Thus, provided we consider both components, so as not to lose any dynamical information, the reduced space is isomorphic to $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the codimension-1 subspace $\widetilde{E}$ may intuitively be identified with a covariant configuration space comprised of all original field variables, except $\rho$. The reduced space does not simply inherit the structure of the jet bundle $J^{1}\widetilde{E}$, since the field variable $\rho$ has been excised, but the velocities $\rho_{\mu}$ have not; indeed these $d$ independent parameters span a copy of $\mathbb{R}^{d}$, and the Cartan form $\Theta_{L^{H}}$ determines whether this defines a Reeb distribution.
Supposing that $(\mathcal{S}_{0},\Theta_{L^{H}})$ does constitute a pre-multicontact system, with characteristic distribution $\mathcal{C}$, we calculate the $(d+1)$-form $\Omega_{L^{H}}:=\bar{d}\Theta_{L^{H}}$, and introduce a multivector field $\boldsymbol{X}_{L^{H}}\in\mathfrak{X}^{d}(\mathcal{S}_{0})$, with local decomposition
@@ XII Effective Non-Abelian Gauge Theory
The appearance of scalar fields that directly couple to gauge curvature terms is a phenomenon which arises most notably in string-inspired models [94, 95]. Indeed, the imprint of dilaton fields on low-energy effective actions has speculatively been regarded as a means to offer novel insight into problems such as colour confinement [96, 97]. In this section, we present an example of a non-Abelian gauge theory coupled to a dilaton-like scalar field; our goal is to illustrate how the concepts of previous sections are applied in practice, as well as to discuss the physical interpretation of our results.
@@ XII.1 Geometrical Setting
Let $\pi:P\rightarrow M$ be a principle $G$-bundle over the $d$-dimensional spacetime manifold $M$; on $P$, there exists a unique, Lie algebra valued connection $1$-form $\omega\in\Omega^{1}(P,\mathcal{L}(G))$, which descends to a local form $A:=s^{*}\omega\in\Omega^{1}(U,\mathcal{L}(G))$, in which $s:U\rightarrow P$ is a local trivialising section of $P$, defined on the open subset $U\subset M$. We shall suppose that the spacetime manifold $M$ is equipped with a Lorentzian metric $g$ of signature $(+,-,-,\,\cdots\,)$; this metric is considered to be parametric, and non-variational [19], so that, physically, our theory is defined on a curved background, which we may change, but is not coupled to gravity. The appropriate covariant configuration space for this theory is
@@
where the trace is taken over the indices of $\mathcal{L}(G)$, and $F$ refers to the Lie algebra valued curvature 2-form $F:=dA+\frac{1}{2}[A,A]$.
In what follows, we shall adopt group-theoretic conventions more prevalent in the physics literature, taking the generators $T_{a}$ of the Lie algebra to be Hermitian, with $T_{a},T_{b}]=if_{ab}^{\;\;\;c}\,T_{c}$. Supposing that $G$ is a semi-simple compact Lie group, the Hermiticity of the $T_{a}$ implies that the Killing form is both non-degenerate, and positive semi-definite; in the fundamental representation, we adopt the following normalisation with respect to the trace
@@
We remark that, with the above conventions, we have implicitly assumed that the dilaton transforms trivially under $G$; were $\phi$ to transform in some non-trivial representation $\rho:G\rightarrow GL(V)$, we would instead have to consider the associated bundle $P\times_{\rho}V$, with the result that our covariant configuration space $\mathcal{E}$ would instead be
$\mathcal{E}=C(P)\times_{M}\left(P\times_{\rho}V\right)\times_{M}\textrm{Sym}_{2}^{1,d-1}(M)$
where the trivial bundle $M\times\mathbb{R}$ has been replaced with the vector bundle $P\times_{\rho}V$ associated to $P$ via the representation $\rho$. In the interest of limiting the number of additional complications, we shall content ourselves with a dilaton that transforms trivially. With these considerations, we rewrite (105) as
@@
in which the curvature $F^{a}_{\mu\nu}$ has the following local coordinate expression
$F_{\mu\nu}^{a}=A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}A^{c}_{\nu}$
The degeneracy of this Lagrangian arises as a result of the gauge symmetry; indeed, the Hessian matrix with respect to $A^{a}_{\mu,\nu}$ and $\phi_{\mu}$ has the following entries
@@ XII.2 Contact Reduction
Upon making the change of variables $\phi\mapsto e^{\rho/2}$, the scaling symmetry vector field $\Sigma$ is now simply $\partial_{\rho}$, and the Lagrangian, with the curvature term expanded in full, reads
$\begin{split}L=-\,\frac{e^{\rho}}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f_{de}^{\;\;\;a}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]\\ +e^{\rho}\left(\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)\end{split}$ (110)
Recall from section (XI) that when the Lagrangian is expressed in the form $L=e^{\rho}f(\rho_{\mu},A_{\mu},A_{\mu,\nu})$, the action density is found from $s^{\mu}=\partial f/\partial\rho_{\mu}$, whilst the Herglotz Lagrangian takes the form $L^{H}=f-\rho_{\mu}s^{\mu}$; for the former, we find
@@
In anticipation of the elimination of $\rho_{\mu}$ in favour of the action density, we have expressed the former in terms of the latter. The Herglotz Lagrangian for this system is then
$L^{H}=-\,\frac{1}{4g^{2}}\biggr[g^{\mu\lambda}g^{\nu\kappa}\left(A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f^{\;\;\;a}_{bc}A^{b}_{\mu}A^{c}_{\nu}\right)\left(A^{a}_{\kappa,\lambda}-A^{a}_{\lambda,\kappa}+f^{\;\;\;a}_{de}A^{d}_{\lambda}A^{e}_{\kappa}\right)\biggr]-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}$ (112)
Due to the singular nature of the Lagrangian (107) from which $L^{H}$ is constructed, it is not a priori guaranteed that the $d$-form $\Theta_{L^{H}}$ endows the reduced space with a pre-multicontact structure. Consequently, prior to implementing the constraint algorithm, we must compute the characteristic and Reeb distributions, verifying that they do indeed satisfy the conditions required for $\Theta_{L^{H}}$ to be pre-multicontact.
Geometrically, referring to the decomposition (103), we see that the excision of the dilaton corresponds to the removal of the trivial bundle $M\times\mathbb{R}\rightarrow M$, defining a reduced covariant configuration space $\mathcal{E}_{\textrm{red}}$. More formally, since $\phi$ is a globally-defined scaling function, the symplectification $\widetilde{\mathcal{E}}_{\textrm{red}}$ is a trivial $\mathbb{R}^{\times}$-bundle over $M$, composed of two connected components $\widetilde{\mathcal{E}}_{\textrm{red}}^{\pm}$, which, with the change of variables $\phi\rightarrow e^{\rho/2}$, correspond to symplectification via $\pm e^{\rho/2}$. Consequently, with the understanding that care must be taken to ensure that both components are considered, we may somewhat informally speak of a reduced space $\mathcal{E}_{\textrm{red}}$ that is isomorphic to $C(P)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$. The Herglotz Lagrangian is defined on the space
@@ XII.3 The Constraint Algorithm
Note that the final quantity on the right-hand side is precisely $L^{H}$, and the bracket $\left(A^{a}_{\nu,\mu}-C^{a}_{\mu\nu}\right)$ vanishes upon imposing semi-holonomy.
The contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ is somewhat involved, and the lengthy algebraic manipulations required are not particularly illuminating; for completeness, we provide the details in appendix A. Performing the contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$, and setting the result to zero, we obtain the following pair of expressions
@@
The first two terms of the second expression form the familiar covariant derivative $(\mathcal{D}_{\mu}F^{\mu\nu})^{a}$, whilst the final term is a common feature of contact-reduced theories, that arises as a result of eliminating a scaling degree of freedom from our ontology. More concretely, in the original Lagrangian (105), there was a coupling between the scalar field $\phi$ and the gauge-kinetic term $\textrm{Tr}\,(F_{\mu\nu}F^{\mu\nu})$; the removal of the dilaton, in favour of a minimal ontology, comes at the cost of introducing a dissipative characteristic into the theory, which is made manifest through the coupling of the gauge field strength to the action density.
@@
Before concluding, there remains an important matter to be discussed, concerning the physical interpretation of the excision of the scalar field $\phi$. In particular, one may be inclined to object that, for theories in which parameter values are determined by expectation values of scalar fields, it is not possible to claim that such fields are ‘redundant’. However, this is not the case; consider, for instance, an experiment in which one attempts to deduce the expectation value $\langle\phi\rangle$, which, for the current discussion, we shall suppose to fix some coupling parameter $g_{s}$. The value of $g_{s}$ is not something directly accessible to an observer, but rather is deduced via comparison to a reference object; this is entirely analogous to how an object of length 1 m is only known to be such because its size is in a 1:1 ratio with that of a previously-standardised instrument. Consequently, under changes generated by $\phi$, the measuring apparatus is subjected to precisely the same rescaling as $g_{s}$, with the result that the overall scaling is empirically inaccessible. Thus, we are justified in asserting that this degree of freedom is redundant, from the perspective of the algebra of dynamical observables.
In fact, we make a somewhat bolder claim: if one is able to construct a theory whose empirically-accessible degrees of freedom are invariant under a particular transformation, we assert that it is a necessary condition, for any relational description, that the mathematical artifice generating such transformations be excisable.
@@
To appreciate why this must be so, consider an arbitrary system $\mathcal{S}$, together with an intrinsic observer, who wishes to construct a relational description of $\mathcal{S}$. The observer has at their disposal only those quantities whose values are ascertainable through physical measurement. For instance, all lengths are deduced with respect to a standard metre stick, and similarly, all times are measured as multiples of the standardised ‘tick’ of a clock.
@@
Suppose that we ourselves are extrinsic ($i.e$ external) observers to $\mathcal{S}$, and we note that, under a certain transformation $\mathcal{T}$, the set of variables we use to describe $\mathcal{S}$ suffer a rescaling. However, this rescaling is such that the dynamical equations we deduce remain invariant. As a trivial example, we consider a pair of variables $(x,t)$, with $\mathcal{T}:(x,t)\mapsto(\lambda x,\lambda t)$; the velocity $\dot{x}$ remains invariant, even though the variables used to deduce $\dot{x}$ do not. The generator of the transformation $\mathcal{T}$ must be a redundant component of our description; if it were redundant, but non-excisable, then it would necessarily appear in the ontology of the observer, else the intrinsic and extrinsic theories would not describe the same system. This contradicts our assumption that the observer’s description is relational. Similarly, if the generator of $\mathcal{T}$ were non-excisable because its presence was required to deduce dynamical evolution, it would be false to claim that the relational description is invariant under $\mathcal{T}$. Consequently, in order for it to be simultaneously true that the observer’s description is wholly relational and that there exists a transformation which leaves the dynamical observables invariant, it must be the case that the generator of this transformation is a feature of our mathematical description that can be removed.

### v2

@@ X.2 The Pre-multicontact Constraint Algorithm
which is clearly an element of $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$. Upon substituting this into the dynamical equations $\iota_{\boldsymbol{X}_{L}}\Theta_{L}=0$ and $\iota_{\boldsymbol{X}_{L}}\Omega_{L}=0$, we must ensure that the resulting expressions do not contain inconsistencies. Supposing that we obtain a compatible set of equations, we now restrict $\textrm{ker}^{d}_{\beta}\,(\Theta_{L},\Omega_{L})$ to the subset of semi-holonomic solutions, which amounts to setting $F_{\mu}^{a}=y^{a}_{\mu}$ in the decomposition (84); in general, this step gives rise to further consistency conditions, and possibly to new constraints.
@@ XI Contact Reduction of Singular Field Theories
Following closely the framework developed in [29], we now present a field-theoretic generalisation of the ideas of section (VII). Let us consider the pre-multisymplectic system $(J^{1}E,\Omega_{L})$, with corresponding Lagrangian function $L:J^{1}E\rightarrow\mathbb{R}$. From the Cartan $d$-form $\Theta_{L}$, we define
@@
The vector field $\Sigma\in\mathfrak{X}^{\infty}(J^{1}E)$ is said to constitute a scaling symmetry of degree $\Lambda$ if
$\mathfrak{L}_{\Sigma}L=\Lambda L\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{\Sigma}\theta_{L}^{\mu}=\theta_{L}^{\mu}\quad\textrm{for all}\;\mu=0,\,\cdots,d-1$ (86)
@@
It is straightforward to verify that the equations of motion derived from $L$ directly imply the Herglotz-Lagrange field equations for $L^{H}$, when the former is restricted to the subspace upon which $L^{H}$ is defined. In light of this, suppose that, having identified a scaling symmetry $\Sigma$ of our pre-multisymplectic system, we adopt coordinates on $J^{1}E$ in such a way so as to render this vector field of the form
@@
for some function $f$. The Euler-Lagrange equation for $\rho$ then implies that $f$ may be written as
@@
which, when compared to (87), and recalling that $\partial_{\mu}s^{\mu}=L^{H}$, suggests we should identify
@@
All steps outlined above are applicable to regular systems; the novelty that arises in the singular case, is that the $d$-form $\Theta_{L^{H}}$ calculated from the symmetry-reduced function $L^{H}$ is not guaranteed to be pre-multicontact. Consequently, we must compute the characteristic and Reeb distributions, after making the ‘naïve’ contact reduction; if these do not satisfy the conditions given in section (X), the symmetry reduction procedure fails. Of course, we shall only consider those physically interesting cases, in which the reduced space does inherit a pre-multicontact structure; nevertheless, it is a non-trivial condition that must be verified.
Having calculated $\Theta_{L^{H}}$, and confirmed that $\textrm{rank}\,\mathcal{D}^{R}=d+k$ and $\textrm{rank}\,\mathcal{C}=k$, for some $0<k\leqslant n(1+d)$, we have a pre-multicontact manifold $(\mathcal{S}_{0},\Theta_{L^{H}})$, in which $\mathcal{S}_{0}$ denotes the reduced space, obtained upon excising the scaling variable $\rho$. Geometrically, if our system is such that the quantity $\xi$ in (89) corresponds to a single field variable on $E$, then, upon writing the Lagrangian in the form (90), $E$ separates into two connected pieces $E_{\pm}$, where each of $E_{\pm}$ is the product of a trivial bundle, and a codimension-1 subspace: $E_{\pm}\cong\widetilde{E}\times_{M}(M\times\mathbb{R}_{\pm})$. The change of variable $\xi\rightarrow e^{\rho/\Lambda}$ selects only the positive component of $\xi$; in order to cover the full range of $\xi$, we must also consider a change of variables in which $\xi\rightarrow-\,e^{\rho/\Lambda}$. Within each of $E_{\pm}$, it is precisely the trivial bundle that is eliminated by the contact reduction. Thus, provided we consider both components, so as not to lose any dynamical information, the reduced space is isomorphic to $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the codimension-1 subspace $\widetilde{E}$ may be identified with a configuration space comprised of all original field variables, except $\rho$. The reduced space does not simply inherit the structure of the jet bundle $J^{1}\widetilde{E}$, as the velocities $\rho_{\mu}$ have not been eliminated; indeed these coordinates span a copy of $\mathbb{R}^{d}$, and the Cartan form $\Theta_{L^{H}}$ determines whether this defines a Reeb distribution.
Supposing that $(\mathcal{S}_{0},\Theta_{L^{H}})$ does constitute a pre-multicontact system, we calculate the $(d+1)$-form $\Omega_{L^{H}}$, and introduce a multivector field $\boldsymbol{X}_{L^{H}}\in\mathfrak{X}^{d}(\mathcal{S}_{0})$, with local decomposition
@@ XII Effective Non-Abelian Gauge Theory
The appearance of scalar fields that directly couple to gauge curvature terms is a phenomenon which arises most notably in string-inspired models [84, 85]. Indeed, the imprint of dilaton-like fields on low-energy effective actions has speculatively been regarded as a means to offer novel insight into problems such as colour confinement [86, 87]. In this section, we present an example of a non-Abelian gauge theory coupled to a scalar field. Having implemented the reduction procedure, we discuss the physical implications of our results.
@@ XII.1 Geometrical Setting
We shall suppose that the spacetime manifold $M$ is equipped with a Lorentzian metric $g$ of signature $(+,-,-,\,\cdots\,)$; this metric is considered to be parametric, and non-variational [19], so that, physically, our theory is defined on a curved background, which we may change, but is not coupled to gravity. The appropriate covariant configuration space for this theory is
@@
where the trace is taken over the indices of $\mathcal{L}(G)$, and $F$ refers to the Lie algebra valued curvature 2-form. In what follows, we shall adopt group-theoretic conventions more prevalent in the physics literature, taking the generators $T_{a}$ of the Lie algebra to be Hermitian, with $T_{a},T_{b}]=if_{ab}^{\;\;\;c}\,T_{c}$. In the fundamental representation, we adopt the following normalisation with respect to the trace
@@
Note that we have implicitly assumed that the dilaton transforms trivially under $G$; were $\phi$ to transform in some non-trivial representation $\rho:G\rightarrow GL(V)$, the trivial bundle $M\times\mathbb{R}$ would be replaced with the vector bundle $P\times_{\rho}V$ associated to $P$ via the representation $\rho$. In the interest of limiting the number of additional complications, we shall content ourselves with a dilaton that transforms trivially. With these considerations, we rewrite (94) as
@@
in which the curvature is expressed locally as $F_{\mu\nu}^{a}=A^{a}_{\nu,\mu}-A^{a}_{\mu,\nu}+f_{bc}^{\;\;\;a}A^{b}_{\mu}A^{c}_{\nu}$. The degeneracy of this Lagrangian arises as a result of the gauge symmetry; indeed, the Hessian matrix with respect to $A^{a}_{\mu,\nu}$ and $\phi_{\mu}$ has the following entries
@@ XII.2 Contact Reduction
Upon making the change of variables $\phi\mapsto e^{\rho/2}$, the scaling symmetry vector field $\Sigma$ is now simply $\partial_{\rho}$, and the Lagrangia reads
$\begin{split}L=-\,\frac{e^{\rho}}{4g^{2}}F_{\mu\nu}^{a}F^{\mu\nu}_{a}+e^{\rho}\left(\frac{1}{8}g^{\mu\nu}\rho_{\mu}\rho_{\nu}-\frac{1}{2}m^{2}\right)\end{split}$ (99)
Recall that when the Lagrangian is expressed in the form $L=e^{\rho}f(\rho_{\mu},A_{\mu},A_{\mu,\nu})$, the action density is found from $s^{\mu}=\partial f/\partial\rho_{\mu}$, whilst the Herglotz Lagrangian takes the form $L^{H}=f-\rho_{\mu}s^{\mu}$; for the former, we find
@@
while the Herglotz Lagrangian for this system is
$L^{H}=-\,\frac{1}{4g^{2}}F_{\mu\nu}^{a}F^{\mu\nu}_{a}-\frac{1}{2}m^{2}-2g_{\mu\nu}s^{\mu}s^{\nu}$ (101)
Geometrically, referring to the decomposition (92), we see that the excision of the dilaton corresponds to the removal of the trivial bundle $M\times\mathbb{R}\rightarrow M$, defining a reduced covariant configuration space $\mathcal{E}_{\textrm{red}}$. More formally, since $\phi$ is a globally-defined scaling function, the symplectification $\widetilde{\mathcal{E}}_{\textrm{red}}$ is a trivial $\mathbb{R}^{\times}$-bundle over $M$, composed of two connected components $\widetilde{\mathcal{E}}_{\textrm{red}}^{\pm}$, which, with the change of variables $\phi\rightarrow e^{\rho/2}$, correspond to symplectification via $\pm e^{\rho/2}$. Provided that both components are considered, we may somewhat informally write $\mathcal{E}_{\textrm{red}}\cong C(P)\times_{M}\textrm{Sym}^{1,d-1}_{2}(M)$. The Herglotz Lagrangian is defined on the space
@@ XII.3 The Constraint Algorithm
Note that the final quantity on the right-hand side is precisely $L^{H}$, and the bracket $\left(A^{a}_{\nu,\mu}-C^{a}_{\mu\nu}\right)$ vanishes upon imposing semi-holonomy. The contraction $\iota_{\boldsymbol{X}}\Omega_{L^{H}}$ produces a pair of expressions, both of which must vanish separately; we provide the details of this calculation in appendix A
@@
Before concluding, there remains an important matter to be discussed, concerning the physical interpretation of the excision of the scalar field $\phi$. In particular, one may be inclined to object that, for theories in which parameter values are determined by expectation values of scalar fields, it is not possible to claim that such fields are ‘redundant’. However, this is not the case; consider, for instance, an experiment in which one attempts to deduce the expectation value $\langle\phi\rangle$, which, for the current discussion, we shall suppose to fix some coupling parameter $g_{s}$. The value of $g_{s}$ is not directly accessible, but must be deduced via comparison to a reference object; this is entirely analogous to how an object of length 1 m is only known to be such because its size is in a 1:1 ratio with that of a previously-standardised instrument. Consequently, changes in $\langle\phi\rangle$ are accompanied by simultaneous rescalings of the sensitivity of our apparatus, with which we would like to ascertain the value of $g_{s}$. It thus follows that the object responsible for these empirically-inaccessible changes in $g_{s}$ (i.e the field $\phi$) is wholly redundant, from the perspective of the physical dynamics.
@@
In fact, we make a somewhat bolder claim: if one is able to construct a theory whose empirically-accessible degrees of freedom are invariant under a particular transformation, we assert that it is a necessary condition, for any relational description, that the mathematical artifice generating such transformations be excisable. To appreciate why this must be so, consider an arbitrary system $\mathcal{S}$, together with an intrinsic observer, who wishes to construct a relational description of $\mathcal{S}$. The observer has at their disposal only those quantities whose values are ascertainable through physical measurement. For instance, all lengths are deduced with respect to a standard metre stick, and similarly, all times are measured as multiples of the standardised ‘tick’ of a clock.
@@
Suppose that we ourselves are external observers to $\mathcal{S}$, and we note that, under a certain transformation $\mathcal{T}$, the set of variables we use to describe $\mathcal{S}$ suffer a rescaling. However, this rescaling is such that the dynamical equations we deduce remain invariant. As a trivial example, we consider a pair of variables $(x,t)$, with $\mathcal{T}:(x,t)\mapsto(\lambda x,\lambda t)$; the velocity $\dot{x}$ remains invariant, even though the variables used to deduce $\dot{x}$ do not. The generator of the transformation $\mathcal{T}$ is clearly a redundant part of our description, and must therefore be excisable. If this were not the case, then it would necessarily appear in the ontology of the observer, else the intrinsic and extrinsic theories would not describe the same system. This contradicts our assumption that the observer’s description is relational. Similarly, if the generator of $\mathcal{T}$ were non-excisable because its presence was required to deduce dynamical evolution, it would be false to claim that the relational description is invariant under $\mathcal{T}$. Consequently, in order for it to be simultaneously true that the observer’s description is wholly relational and that there exists a transformation which leaves the dynamical observables invariant, it must be the case that the generator of this transformation is a feature of our mathematical description that can be removed.